OpenStudy (anonymous):

solve system by any method elimination, substitution . {x+y+3z=-4 -x-y-2z=5 2x-z=-3}

OpenStudy (nurali):

x+y+3z=-4----(1) -x-y-2z=5----(2) 2x-z=-3------(3) in equation 3 2x-z=-3 2x=-3+z 2x=z-3 x=z-3/2-----(4) put x in equation 2 -x-y-2z=5 -(z-3/2)-y-2z=5 3-z/2-y-2z=5 Taking L.C.M 3-z-2y-4z=10 3-5z-2y=10 2y=3-10-5z 2y=-7-5z y=-7-5z/2----(5) put x and y in equation 1 x+y+3z=-4 z-3/2-7-5z/2+3z=-4 Taking L.C.M z-3-7-5z+6z=-8 -4z+6z-10=-8 2z=-8+10 2z=2 z=1 put z in equation 4 x=z-3/2 x=1-3/2 x=-1 put z in equation 5 y=-7-5z/2 y=-7-5/2 y=-12/2 y=-6 {x,y,z}={-1,-6,1}